OpenStudy (emma-.-):

What is the angle of 135 degrees equal to in radian measure?

OpenStudy (marigirl):

\[135 * (\frac{ \pi }{ 180 })= \frac{ 135\pi }{ 180 }=\frac{ 3\pi }{ 4 }\]

OpenStudy (emma-.-):

Thank you